A rotary type compressor compresses air from 98kPa, 27C through a compression ratio of 5 and compression efficiency of 75%, find the compressor work.
Answer #1

Given data:

P_1=98\:kPa,\:T_1=300\:K,\:r=5,\:\eta =0.75

The compression ratio

r=\frac{P_2}{P_1}=5

\therefore P_2=5\times P_1=5\times 98=490\:kPa

Specific volume,

\vartheta_1 =\frac{1}{\rho_1}= \frac{RT_1}{P_1}=\frac{287\times300}{98\times10^3}=0.87857\:\frac{m^3}{kg}

The actual work transferred to air

W_a=\vartheta_1 (P_2-P_1)

W_a=0.87857 (490\times10^3-98\times10^3)=344400\:J/kg

The work consumed by the compressor

W=\frac{W_a}{\eta }=\frac{344400}{0.75}=459200\:\frac{J}{kg}

\therefore\:W=459.2\:\frac{kJ}{kg}